OpenStudy (anonymous):

a jet plane traveling at 550 mph overtakes a propeller plane traveling at 200 mph that had a 4-hour head start. How far from the starting point are the planes?

OpenStudy (radar):

Use algebra with the fact the Distance (d) = Rate (r) times time (t) and further you do know that the distance traveled by each plane was equal at the point where jet overtook the prop plane.

OpenStudy (radar):

Let t = time of Jet plane flight at the point of overtake. Let t + 4 = time of prop plane flight (from start to overtake point) The distance for each plane is equal, and furthermore we know that d=rt. Step 1. calculate distance of Jet plane. d=rt, so d=550 t Step 2. calculate distance of prop pland d=rt, so d=200(t + 4) = 200t + 800 Since the distances are equal, Step 3. Set up the equation: 550t = 200t + 800 Now solve for t and calculate the distance which is the answer that is being sought. d=rt Use the Jets rate and time. and then verify by solving for distance of the prop (using t + 4) Both distances should be equal verifying the correctness of your answer.